OpenStudy (anonymous):

I don't know how to solve indefinite integral of [x^2 * sqrt(3-x)].

OpenStudy (anonymous):

\[\int\limits x^2 \left ( -x+3 \right )^{\frac{1}{2}} dx\]

OpenStudy (dumbcow):

substitution u = 3-x --> x = 3-u du = -dx \[= -\int\limits (3-u)^{2} \sqrt{u} du\] \[= -\int\limits \sqrt{u}(u^{2}-6u+9) du\] distribute and then integrate each term using power rule

OpenStudy (anonymous):

thank you

OpenStudy (dumbcow):

yw

OpenStudy (anonymous):

\[\int\limits x^2 \sqrt{3-x}\:dx = -\frac{2}{7}(3-x)^\frac{7}{2} + \frac{12}{5}(3-x)^\frac{5}{2} - 6(3-x)^\frac{3}{2} + C\]

OpenStudy (dumbcow):

the coefficients seem to be a little off http://www.wolframalpha.com/input/?i=integral+x%5E2+sqrt%283-x%29+dx